A mathematics professor assigns two problems for homework and knows that the probability of a student solving the first problem is 0.75, the probability of solving the second is 0.50, and the probability of solving both is 0.25. (Round your answers to three decimal places.) (a) Jed has solved the second problem. What is the probability he also solves the first problem? (b) Edna has solved the first problem. What is the probability she also solves the second problem?
Answer #1

p( solung ist ) = 0.45 = P(A) pcsolung 2nd) = 0.50 = P(B) pcsolung berk) = 0.25 - P(ANB) a PC solves ist | solved 2nd) EPCA/B